OpenStudy (anonymous):

simplify sqrt 5- sqrt 10 ------------ sqrt 5+sqrt 10

OpenStudy (angela210793):

sqrt 5- sqrt 10 (sqrt 5-sqrt 10)*(sqrt 5-sqrt 10) (sqrt 5-sqrt 10)^2 ---------------=--------------------------------=------------------- (sqrt 5+sqrt 10) (sqrt 5+sqrt 10)*(sqrt 5-sqrt 10) 5-10 (sqrt 5-sqrt 10)^2 =----------------- -5

OpenStudy (anonymous):

\[(\sqrt{5}-\sqrt{10})/(\sqrt{5}+\sqrt{10})\] \[(\sqrt{5}-\sqrt{10})*(\sqrt{5}-\sqrt{10})\div(\sqrt{5}+\sqrt{10})(\sqrt{5}-\sqrt{10})\] \[5+10-2*(\sqrt{5}*\sqrt{10}) \] --------------------- (5-10)